OpenStudy (brinazarski):

Which polar coordinates represent the same point as (3, π/3)? (Select all that apply.)

OpenStudy (brinazarski):

(−3, −2π/3) (−3, 4π/3) (3, 5π/3) (3, −π/3) (−3, −7π/3) (3, 7π/3)

OpenStudy (brinazarski):

I only selected the last one but there's more

OpenStudy (anonymous):

have you tried drawing it out?

OpenStudy (brinazarski):

Would it be the last 3 only?

OpenStudy (anonymous):

if you take the time to draw out each of the polar coordinates, then you should have your answer

OpenStudy (brinazarski):

But they're all right triangles...

OpenStudy (anonymous):

Where on the polar coordinate graphs do those points lie? They're not right triangles at all, they are just points on a graph, single 0-dimensional points. Just like on a cartesian coordinates graph, there are points, so are points on a polar coordinates graph

OpenStudy (jdoe0001):

keep in mind that \(\bf \cfrac{7\pi}{3}\implies \cfrac{6\pi}{3}+\cfrac{\pi}{3}\implies 2\pi+\cfrac{\pi}{3}\)

OpenStudy (brinazarski):

I don't know much about polar coordinates so...

OpenStudy (ranga):

\[(r, \theta \pm 2n \pi) \quad \text{and} \quad (-r, \theta \pm (2n+1) \pi)\] will all represent the same point: (r, theta) where n = 0, 1, 2, 3, .... (integer)

OpenStudy (brinazarski):

@jdoe0001 so the last one should work, right?

OpenStudy (anonymous):

Cartesian Coordinates (X-axis point, Y-axis point) Polar Coordinates (Radius from origin point, How many units traveled in Radians(Unit Circle) point)

OpenStudy (brinazarski):

But the second to last one shouldn't work...?

OpenStudy (ranga):

The second one will work. Putting a negative in front of r and adding pi to the angle will represent the same point.

OpenStudy (brinazarski):

So would the first one work?

OpenStudy (ranga):

The first will work too.

OpenStudy (brinazarski):

So when r is positive, add/subtract 2pi, when r is negative, add/subtract pi?

OpenStudy (ranga):

exactly.

OpenStudy (ranga):

multiples of 2pi and multiples of pi

OpenStudy (brinazarski):

So it's only the first two, the fourth, and the last one?

OpenStudy (ranga):

or you can simply draw each one out as suggested before.

OpenStudy (ranga):

pi/3 = 60 degrees 2pi/3 = 120 degrees, etc.

OpenStudy (ranga):

-r is same as +r in the opposite direction.

OpenStudy (brinazarski):

Sooo... those aren't the answers? I'm still confused cuz I thought those were the answers based on this but it's still not accepting and I have one try left. The remaining ones dont look like answers...

OpenStudy (jdoe0001):